OpenStudy (anonymous):

using differentiation, how do you solve this? find two positive numbers whose product is 64, and whose sum is a minimum. (the answer should be: (8,8) )

OpenStudy (anonymous):

Let the two positive numbers be x and x/64 Now form y=x+x/64 Take derivative of y with respect to x Now carry on...........

OpenStudy (anonymous):

why is the other number x/64?

OpenStudy (anonymous):

because the product of x and x/64 is 64. Check it out by yourself

OpenStudy (anonymous):

Wouldn't it be x and 64/x

OpenStudy (anonymous):

Sorry yes x and 64/x

OpenStudy (anonymous):

What a shame!

OpenStudy (anonymous):

then, why not 64/x?

OpenStudy (anonymous):

Now y=x+64/x And differentiate

OpenStudy (anonymous):

Sorry! My blunder, What the hell!

OpenStudy (anonymous):

it's alright.

OpenStudy (anonymous):

Now can you work out from here

OpenStudy (anonymous):

so, it will be 1 + (-64/x^2) =y when differentiated?

OpenStudy (anonymous):

Yes

OpenStudy (anonymous):

Yeap, now equate it with 0

OpenStudy (anonymous):

You will get x=8 and -8

OpenStudy (anonymous):

x=8 and x =-8 are they the answers?

OpenStudy (anonymous):

but the number should be positive so choose 8

OpenStudy (anonymous):

Hence the numbers are 8 and 64/8=8

OpenStudy (anonymous):

because a minimum is > 0? is that it?

OpenStudy (anonymous):

A maximum or minimum is when the first derivative is equal to zero.

OpenStudy (anonymous):

Second derivative will be greater than zero check it out by yourself

OpenStudy (anonymous):

Also the first derivative is always zero for minimum and maximum

myininaya (myininaya):

\[xy=64\] <-two positive numbers whose product is 64 Let S be the sum of x and y \[\ S=x+y\] so we know xy=64 => y=64/x so we can write S using one variable \[\ S=x+\frac{64}{x}\] now S' needs to be found since we are trying to minimize \[\ S'=1-\frac{64}{x^2}\] now we need to set S'=0 \[\frac{x^2-64}{x^2}=0\] 8,-8, and 0 are critical numbers but we throw out -8 and 0 since x>0 (remember we looking for numbers and also if x=0 then the product would not be 64 it would be 0) so we have x=8 => y=64/8=8

myininaya (myininaya):

positive numbers*